OpenStudy (anonymous):

Fill in the blank to make the expression a perfect square. y^2-6y+?

OpenStudy (anonymous):

+9

OpenStudy (anonymous):

:D let me know if u have any questions

OpenStudy (anonymous):

thanks! :)

OpenStudy (anonymous):

wait, how did yu get that

OpenStudy (anonymous):

okay let me explain

OpenStudy (anonymous):

\[y ^{2}-6y+x ^{2} = (y-x)^{2}\]

OpenStudy (anonymous):

simplifying this equation will give you x=3

OpenStudy (anonymous):

so the missing value is 9

OpenStudy (anonymous):

oh okay, I understand now :)

OpenStudy (anonymous):

no problemo

OpenStudy (whpalmer4):

if you have something like \[y^2-6y\] and you want to make it a perfect square, just take half of the coefficient of the \(y\) term, square it, and add that. Then you can rewrite it as \((y+a)^2\) where \(a =\) the value you squared. \[y^2-6y + (\frac{-6}{2})^2 = y^2-6y + 9 = (y-3)^2\] This is called "completing the square" for what should be fairly apparent reasons :-)

OpenStudy (whpalmer4):

If you were solving an equation by this method, you would add the squared quantity to both sides of the equation to preserve the equality. \[x^2+4x =0\]\[x^2+4x+(\frac{4}{2})^2 = 0+(\frac{4}{2})^2\]\[(x+2)^2 = 4\]\[\sqrt{(x+2)^2} =\pm \sqrt{4}\]\[x+2 = \pm2\]\[x=0,x=-4\]

OpenStudy (whpalmer4):

as a check, solve by factoring: \[x^2+4x=0\]\[x(x+4) = 0\]\[x=0\]\[x+4=0\]\[x=-4\]